A channel has trapezoidal ends. The top width of the channel is measured to be 6m and the bottom width is 3m. If the

depth of the channel is 4m. Find the volume per meter strip.


1
Expert's answer
2021-09-22T00:22:38-0400

Area

"A=\frac{a+b}{
h}
\\
A=\frac{3+6}
2·4=18m^2\\
Volume = Area * length \\
Volume = 18m^2*m\\
Volume =18 m^3"